# Out-of-Distribution Runtime Adaptation with Conformalized Neural Network Ensembles

### v1

@@ VII-A Dynamics Model
Figure 4: The components of the state of the Reeds-Shepp Vehicle, where ${x_{t}\in\mathbb{R}^{2}}$ denotes its position, ${V_{t}\in\mathbb{R}}$ its speed, ${\theta_{t}\in\mathbb{R}}$ its heading-angle, and ${\kappa_{t}\in\mathbb{R}}$ the curvature of its trajectory.
VII-B Dataset
We use the the Vehicle-Crowd Intraction[sic] (VCI) dataset [47], to train and evaluate the performance of each method in simulation. The dataset used, a subset of the total VCI dataset, consists of $110$ trajectories of pedestrians crossing a roadway, with each trajectory comprised of the position data of a pedestrian for $154$ time steps. In our evaluations, we randomly initialize the position of the pedestrian, relative to the ego vehicle. The horizontal starting positions are sampled from a normal distribution with a mean $40$ meters from the horizontal starting position of the ego vehicle and a standard deviation of $2.5$ meters. The starting vertical position of the pedestrian is determined by the direction of travel in the specific trajectory obtained from the dataset. To generate the training data, we process the dataset into input-output data pairs, where the input consists of the positions of the pedestrian over $14$ consecutive time steps and the output consists of the position of the pedestrian at the following time step.
Of the 110 trajectories, 10 are selected randomly to serve as the test dataset $D_{\mathrm{test}}$, used in evaluating the performance of each algorithm. For the remaining 100 trajectories, one data pair from each trajectory is selected uniformly randomly and removed, with these 100 data points forming the calibration dataset $D_{\mathrm{cal}}$ used in conformal prediction. The remaining data points comprise the training dataset $D_{\mathrm{train}}$, used in training the parameters of each neural network in the ensemble through Adam optimization. These neural networks were used in testing for nearly all of the methods and benchmarks, with the exception of the testing for benchmarks based on [22] and [27]. For the benchmark based on [27], no calibration dataset was formed, and all input-output data points from the 100 trajectories were used for training. For the benchmark based on [22], 20 uniformly randomly selected trajectories from the 100 full trajectories were used to form the calibration dataset, with the remaining 80 trajectories used in training the neural networks.
One of the limitations of the VCI dataset is the amount of data available: with only 15400 data points in total available across training, calibration and test datasets, it is not possible to make statistical determinations that depend on large amounts of data. To remedy this, a neural network more sophisticated than those used for the predictor ensemble was trained on the existing trajectories and used to generate additional pedestrian trajectories, used for the calibration process and statistical analysis. The architecture of the network used to generate new data can be found in Figure 10, where the neural network uses an LSTM layer with 128 neurons, a dense layer of 128 neurons and ReLU activations to transform rectilinear trajectories of 2-dimensional states into trajectories of two-dimensional states mimicking the behavior of a pedestrian.
VII-C Predictor Architecture
VII-D OOD Detector
In our simulations, we utilize ${K=97}$, for $N=100$ (i.e. $\delta\approx 3.9604\%$) in calibrating the OOD detector, which yielded good results for reliable OOD detection, with $C$ set to the value of $\rho^{(K)}$ rounded up to the third decimal point. Careful readers will notice that the probabilistic guarantees are specifically placed on the probability of correctly determining that a given data point was generated by the same distribution that generated the training data, but as we will show in the section detailing the results of our experiments, over the course of our experiments we were able to reliably detect changes in the underlying data-generating distribution.
In our simulations, our OOD detector classifies any data point with an uncertainty score above $C$ as being out-of-distribution. We note, however, that the simplicity of this detection framework can result in incorrect classifications of in-distribution samples as OOD, given the probabilistic guarantees in Theorem 1. Nevertheless, we can address this challenge using existing methods such as [4], which enables the detector to ignore occasional high-uncertainty scores from in-distribution points, but still respond appropriately to persistent high-uncertainty scores that indicate a distribution shift. Further, other uncertainty metrics in anomaly detection, such as those in [33] and [19] can be applied in OOD detection. However, these methods require access to the training and calibration data, which could pose an additional challenge, unlike our method.
VII-E Empirical Results
We assess the performance of our control algorithm OOD-MPC against four baselines, namely: (a) SPDE [22]; (b) MARC [27]; (c) Reachable Sets Only, a variant of OOD-MPC utilizing only the MPC II controller; and (d) Ensembles Only, a variant of OOD-MPC utilizing only the MPC I controller. We examine each algorithm in $20$ simulations, presenting the results achieved by each algorithm in the subsequent discussion. Each simulation terminates in one of three ways: (a) the autonomous vehicle goes past the pedestrian while avoiding collisions, (b) the autonomous vehicle does not go past the pedestrian but avoids a collision (e.g., when stopping to allow the pedestrian to finish crossing the road), or (c) the autonomous vehicle collides with the pedestrian. In OOD-MPC, we utilize a failure probability of approximately $4\%$ in the conformal prediction procedure for computing prediction regions, associated with ${C=0.012}$. In contrast, the uncertainty quantification for the MARC algorithm [27] corresponds to a failure probability of $5\%$ and the ball-shaped confidence areas for the SPDE algorithm [22] correspond to a failure probability of $5\%$. In all the figures, we depict the past trajectory of the autonomous vehicle in blue when the autonomous vehicle detects nominal behavior and in red when the autonomous vehicle detects out-of-distribution behavior. Further, we depict the past trajectory of the pedestrian in cyan.
Figure 5 shows the results of the simulations for all algorithms. We note that OOD-MPC and the more conservative methods MARC and Reachable Sets Only do not collide with the pedestrian, even when the pedestrian attempts to force a collision, while exhibiting the insurance fraud behavior, unlike the other algorithms. In the nominal case, the Ensembles Only method achieves the highest success percentage at $70\%$, without any collisions, followed jointly by OOD-MPC and SPDE, which achieve a success percentage of $60\%$. Consequently, OOD-MPC provides a desirable tradeoff, suitable across both behavior modes.
Figure 5: Results of simulations for each algorithm when the pedestrian acts either nominally or attempts to force a collision (insurance fraud). Of the implemented strategies, only OOD-MPC (Our Method) is able to both pass the horizontal position of the pedestrian in the nominal case and successfully avoid collisions with the pedestrian in the insurance fraud case.
We note some important findings associated with each algorithm in the course of the simulations. While the Ensembles Only method could reliably avoid collisions in the nominal case, the method failed to avoid collisions in the insurance fraud case, since the predictions of the pedestrian trajectory were notably different from the actual trajectory taken by the pedestrian, as depicted in Figure 6. In contrast, the Reachable Sets Only method avoids collisions in all settings; however, with this method, the autonomous vehicle was consistently unable to reach its desired goal location during the simulation, as illustrated in Figure 7. We note that the MARC algorithm achieved similar conservative results to those of the Reachable Sets Only method.
Figure 6: Ensembles Only: (Top) In the nominal case, the Ensembles Only method avoids collisions. However, in the insurance fraud setting (Bottom), the algorithm makes incorrect predictions of the future trajectory of the pedestrian, resulting in a collision.
Figure 7: Reachable Sets Only: The Reachable Sets Only method avoids collisions in both the nominal (Top) and insurance fraud (Bottom) behavior settings. However, the autonomous vehicle does not reach its desired goal location.
Although the SPDE algorithm enables the autonomous vehicle to reach its goal location while avoiding collisions in the nominal case, this algorithm does not avoid collisions when the pedestrian attempts to forcefully collide with the vehicle: this can be attributed to the erroneous predictions generated by the neural network ensemble. Though the method adds a buffer around the position of the pedestrian based on the estimated uncertainty of its prediction, the uncertainty is calibrated under the assumption that the true pedestrian trajectory is drawn from the same distribution as in the training data, while in the insurance fraud setting it is drawn from an altogether different distribution. In contrast, the OOD-MPC algorithm was not only able to avoid collisions in the nominal case, but also avoids collisions in all the simulations in the insurance fraud setting, where the controller switches from the nominal MPC I control strategy to the reachable-set-based MPC II control strategy after detecting out-of-distribution behavior, depicted in Figure 8.
Figure 8: OOD-MPC: The OOD-MPC algorithm avoids collisions in both the nominal (Top) and insurance fraud (Bottom) behavior settings. The autonomous agent correctly detects out-of-distribution behavior and switches its control strategy to the reachable-set-based control strategy in the insurance fraud setting.
VII-F Statistical Analysis of Safety Assurances on Real and Synthetic Data
@@
Here, we analyze the statistical guarantees of our out-of-distribution detector empirically. We note that conformal prediction only provides probabilistic guarantees on the false-positive rate (i.e., on an incorrect identification of a data point in a nominal trajectory as an out-of-distribution sample). To assess the tightness of this guarantee, we apply the out-of-distribution detector at $27$ time steps in each of the $10$ nominal simulations, and at the $23$ corresponding time steps after the shift in pedestrian behavior from nominal to OOD behavior in each of the $10$ insurance fraud simulations. We summarize these results in Table I, noting that the OOD detector has a false-positive rate of ${4.4\%}$, which is close to the probabilistic error rate of ${4\%}$, provided by conformal prediction. In Figure 9, we show a simulation where the detector incorrectly classifies a nominal behavior as being OOD. However, our detector is able to recover from incorrect classifications and subsequently classifies the behavior of the pedestrian as being in-distribution.
@@
TABLE I: Confusion matrix showing the performance of the out-of-distribution (OOD) detector on the pedestrian trajectory data. The percentage values are computed relative to the ground-truth nominal and OOD behaviors: nominal cases cover all 10 nominal trials and OOD cases cover all 10 insurance fraud trials after the shift in pedestrian behavior. The calibrated error rate $\delta\approx 4\%$ is close to the empirical false-postive rate associated with nominal data points identified as OOD (shown in bold).
@@
Figure 9: Our OOD detector is able to recover from incorrect classifications. Here, the detector wrongly classifies nominal behavior as being OOD, leading to a temporary switch from the nominal MPC controller (blue trajectory) to the conservative, provably-safe controller (red trajectory), before making the correct classification.
@@
Figure 10: Architecture of the LSTM neural network used to generate synthetic trajectory data for statistical analysis.
We note that the discrepancy between the observed error rate and the statistical guarantee provided by conformal prediction is expected: more specifically, since each of the trajectories are sampled i.i.d. from the probability distribution $\mathcal{D}$, the distribution of coverage for a threshold calibrated with $C=\rho^{(K)}$ follows a Beta distribution, with shape parameters $(a,b)=(N+1-K,K)$ [2].
In Figure 11, we show the empirical distribution of coverage, in addition to the associated Beta distribution. We generate the empirical distribution using synthetic trajectories generated by an LSTM (described in Figure 10), trained on our training dataset, where the trajectories for the calibration and validation datasets are independently sampled without replacement. For each of $3000$ independent trials, $100$ independent trajectories are first generated and used to create a new calibration set of $100$ i.i.d. data points, with a new threshold set equal to the covariance measure associated with ${K=97}$ (the neural networks used to generate the covariance score are the same as during the simulations, and were not retrained for this process). Subsequently, the covariance score for each of the data points in a further $150$ independently sampled trajectories is compared to the newly calibrated threshold, with the empirical coverage achieved during each trial being defined as the fraction of the data points in these $150$ trajectories with a covariance score less than or equal to the calibrated threshold. The number of trials that are associated with a given empirical coverage, out of the $3000$ in total, are proportional to the probability density described by the Beta distribution. Figure 11 shows that the empirical distribution closely matches the Beta distribution.
@@
Figure 11: Beta distribution based on calibration set of 100 samples and a marginal probability of $96\%$, with empirical results from validation sets of $150$ trajectories each showing the results of $3000$ trials with LSTM-generated trajectories. The empirical coverage rate seen in Table I, $95.6\%$, is also listed.
VIII Limitations
Covariance estimation in neural network ensembles requires the outputs of the individual models to differ. We note that this requirement precludes the utilization of linear predictors trained to the global optimum associated with this architecture. Further, our proposed algorithm does not distinguish between different categories of out-of-distribution behavior of dynamic agents.
IX Conclusion
We introduce a safe, adaptive controller that enables autonomous robots to safely navigate in their environments in the presence of other dynamic agents, with nominal or out-of-distribution behavior. Our adaptive controller is comprised of a nominal control strategy based on neural network predictions for in-distribution settings, and a more conservative control strategy based on reachable sets for out-of-distribution cases. Further, we introduce an out-of-distribution detector, which utilizes conformal prediction to provide provable probabilistic guarantees on the validity of the prediction regions. We demonstrate the safety and efficacy of our controller in simulation, where an autonomous vehicle attempts to reach a goal location without colliding with a pedestrian crossing the road at a crosswalk. With our algorithm, the autonomous vehicle avoids collisions with the pedestrian in all settings, and is able to pass the position of the pedestrian when the pedestrian acts in-distribution; in the same scenario, some other existing methods either result in collisions, when the pedestrian attempts to force them, or are overly conservative (failing to reach the goal).
In future work, we seek to explore the utilization of deep-learned controllers in combination with (or in lieu of) MPC-based controllers in our control framework, which could be particularly advantageous in problems involving robots with more complex dynamics models. Such work could enable the development of more expressible control frameworks. In addition, we are interested in exploring the use of our method in uncertainty quantification to identify and search for areas of the predictor input space not appropriately covered by training data, and in examining other uncertainty quantification metrics for out-of-distribution detection.
X Appendix
X-A Trajectory Prediction
To generate predictions, at each time step, for the movements of a pedestrian for the time remaining in the simulation, we recursively apply predictions for the number of time steps remaining in the simulation. After making a prediction at time step $t$ of the pedestrian’s position at time step $t+1$ (using the pedestrian’s positions during the preceding $\ell$ time steps), we query the network ensemble again and define the input to the network by combining the positions corresponding to the preceding $\ell-1$ time steps and the prediction made for $t+1$ in order to get a prediction for the pedestrian position at $t+2$; we subsequently make predictions further into the future in this way until we reach the number of remaining time steps in the simulation. To predict pedestrian movements for the initial time steps (before enough time steps have elapsed in a simulation for the network ensemble to make predictions), we assume that the pedestrian moves at $1.1m/s$, a speed close to the mean pedestrian speed recorded in the trajectories of the training set, with the direction of movement given by the starting position of the pedestrian.
This particular neural network design, and implementation of multi-step predictions, was selected to minimize the complexity of the neural network architecture. While the method of prediction is less sophisticated and computationally efficient than many state-of-the-art implementations, the reduced complexity allows us to emphasize that the principles of neural networks exploited to detect out-of-distribution behavior work in small-scale implementations, and are not emergent behavior of more sophisticated models.
X-B Model Predictive Control
The MPC optimizer uses an implementation of sequential convex programming, including affine approximations of nonconvex constraints. The optimization problem is run multiple times, until convergence of the solution (initial trajectory estimates, for the very first iteration of the convex approximate problem, depend on the control strategy: in the case of the control strategy based on neural network predictions, they first consist of linear interpolations of the starting and goal states, and then of a trajectory that avoids obstacles but ignores constraints imposed by control limits; in the case of the conservative control strategy, the initial trajectory consists of repeated instances of the starting state). Once convergence is reached for a given solution, the MPC optimizer computes full, distinct trajectories of states and control inputs to reach as close to the target goal as possible. In this particular implementation, the solver resolves the problem every $H=5$ time steps, each time calculating the trajectory for the vehicle to follow for the time remaining in the simulation.
X-C Simulation Setup
The simulation environment is based on a two-lane road, $7.2$ meters wide in total. The goal of the ego vehicle is to move from an initial position, on the middle of the right lane, with an initial velocity of $10\mathrm{m/s}$ parallel to the road, to a goal position in the middle of the right lane $70\mathrm{m}$ further along the road, at the same velocity. In the objective function, we penalize the error between the vehicle’s position and the goal position more than the error between its velocity and the goal velocity. Each simulation runs for $150$ time steps at $23.976\mathrm{Hz}$, the frequency at which the original pedestrian data was sampled, meaning each simulation represents approximately $6.256$ seconds in real time.
The dimensions of the autonomous vehicle are based on the dimensions of the 2019 Ford Fiesta [6], with a width of $1.8$ meters, a length of $4.0$ meters and a minimum turn radius of approximately $5.913$ meters. We assume the autonomous vehicle has a maximum speed of $20\mathrm{m/s}$, while the pedestrian has a maximum speed of $4.5\mathrm{m/s}$, close to the speed at which an average person can run for short periods of time. The vehicle has two radii for collision and obstacle avoidance: $0.9\mathrm{m}$, based on vehicle width, which is the minimum distance that the vehicle must keep from the upper and lower road boundaries, and $2.0\mathrm{m}$, based on vehicle length, which is the minimum distance that the vehicle must keep from the pedestrian to prevent a collision. The pedestrian is represented as a circular object with a radius of $0.5\mathrm{m}$, and is assumed to be able to move in any direction on a planar surface, with the state at time step $t$, $\mathbf{s}_{t}\in\mathbb{R}^{2}$, consisting of the position of the pedestrian in two dimensional space. In these simulations, we consider a single pedestrian interacting with an autonomous vehicle. However, we note that our proposed algorithm applies to problems with multiple pedestrians, or problems with both pedestrians and other vehicles.
X-D Synthetic Data
To train the neural network used in generating synthetic data, original pedestrian states were used as example outputs: $\mathbf{s}_{1:T}$ represents all of the states in one trajectory with $T$ time steps. Simultaneously, the corresponding inputs to each trajectory, $\mathbf{r}_{1:T}$, were generated by taking the starting and ending states of the trajectory such that $\mathbf{r}_{1}:=\mathbf{s}_{1}$ and $\mathbf{r}_{T}:=\mathbf{s}_{T}$, then finding linear interpolations of these states. New, synthetic trajectories were generated after training by creating new rectilinear trajectories to serve as inputs to the network: as average pedestrian speeds and the angles between their starting and ending states each approximated a normal distribution, new rectilinear trajectories were created by sampling from normal distributions following the parameters matching those of the real data and passing the resulting trajectories through the network. The vertical direction then had a 50% chance of being inverted, and the horizontal starting position of the pedestrian relative to the starting position of the ego vehicle was normally distributed as the trajectories drawn from the VCI dataset were.
X-E Pedestrian Behavior
For each simulation, the pedestrian adopts one of two modes of behavior for the whole trajectory. For each of the $10$ trajectories in the test set, the pedestrian either follows the selected trajectory throughout while crossing the crosswalk (representing in-distribution behavior for the predictor ensemble), referred to as nominal behavior, or the pedestrian can follow one of these trajectories for a set period of time (in these simulations, $1.3$ seconds) and then suddenly change its behavior by continuously moving towards the ego vehicle at maximum speed, referred to as insurance fraud behavior, representing a pedestrian attempting to force a collision with the ego vehicle. This presents both a worst-case scenario for the ego vehicle, which must prevent a collision despite the restrictions on its movements, and a shift in distribution that the OOD detector must identify and respond to over the course of the simulation.
X-F Statistical Analysis for Data Collection

### v3

@@ C.4 Dynamics Model
Figure 5: State of the Reeds-Shepp Vehicle, with its position ${x_{t}\in\mathbb{R}^{2}}$, its speed ${V_{t}\in\mathbb{R}}$, its heading-angle ${\theta_{t}\in\mathbb{R}}$, and curvature ${\kappa_{t}\in\mathbb{R}}$.
C.5 Trajectory Prediction
To generate predictions at each time step for the movements of a pedestrian for the time remaining in the simulation, we recursively apply predictions for the number of time steps remaining in the simulation. Figure 6 shows the architecture of each predictor. We note that the individual models comprising the ensemble are not required to have the same architecture. After making a prediction at time step $t$ of the pedestrian’s position at time step $t+1$ (using the pedestrian’s positions during the preceding $\ell$ time steps), we query the network ensemble again and define the input to the network by combining the positions corresponding to the preceding $\ell-1$ time steps and the prediction made for $t+1$ in order to get a prediction for the pedestrian position at $t+2$; we subsequently make predictions further into the future in this way until we reach the number of remaining time steps in the simulation. To predict pedestrian movements for the initial time steps (before enough time steps have elapsed in a simulation for the network ensemble to make predictions), we assume that the pedestrian moves at $1.1m/s$, a speed close to the mean pedestrian speed recorded in the trajectories of the training set, with the direction of movement given by the starting position of the pedestrian.
This particular neural network design and implementation of multi-step predictions was selected to minimize the complexity of the neural network architecture. While the method of prediction is less sophisticated and computationally efficient than many state-of-the-art implementations, the reduced complexity allows us to emphasize that the principles of neural networks exploited to detect out-of-distribution behavior work in small-scale implementations and are not emergent behavior of more sophisticated models.
C.6 Model Predictive Control
The MPC optimizer uses an implementation of sequential convex programming, including affine approximations of nonconvex constraints. Starting with an initial estimate for a trajectory that does not incorporate physical constraints, the optimization problem is run multiple times until the solution converges. The initial estimate used depends on the control strategy. When the control strategy is based on neural network predictions, the estimate first consists of linear interpolations of the starting and goal states, and then, of a trajectory that avoids obstacles but ignores constraints imposed by control limits. For the conservative control strategy, the initial estimate consists of repeated instances of the starting state. Once an initial estimate is produced, the MPC optimizer computes full, distinct trajectories of states and control inputs to reach as close to the target goal as possible by iteratively modifying the solution until convergence is reached. In this particular implementation, the solver resolves the problem every $H=5$ time steps, each time calculating the trajectory for the vehicle to follow for the time remaining in the simulation.
C.7 Simulation Setup
The simulation environment is based on a two-lane road, $7.2$ meters wide in total. The goal of the ego vehicle is to move with an initial velocity of $10\mathrm{m/s}$ parallel to the road from an initial position in the middle of the right lane to a goal position in the middle of the right lane $70\mathrm{m}$ further along the road at the same velocity. In the objective function, we penalize the error between the vehicle’s position and the goal position more than the error between its velocity and the goal velocity. Each simulation runs for $150$ time steps at $23.976\mathrm{Hz}$, the frequency at which the original pedestrian data was sampled, meaning each simulation represents approximately $6.256$ seconds in real time.
The dimensions of the autonomous vehicle are based on the dimensions of the 2019 Ford Fiesta (Company, 2018), with a width of $1.8$ meters, a length of $4.0$ meters and a minimum turn radius of approximately $5.913$ meters. We assume the autonomous vehicle has a maximum speed of $20\mathrm{m/s}$, while the pedestrian has a maximum speed of $4.5\mathrm{m/s}$, close to the speed at which an average person can run for short periods of time. The vehicle has two radii for collision and obstacle avoidance: $0.9\mathrm{m}$, based on vehicle width, which is the minimum distance that the vehicle must keep from the upper and lower road boundaries, and $2.0\mathrm{m}$, based on vehicle length, which is the minimum distance that the vehicle must keep from the pedestrian to prevent a collision. The pedestrian is represented as a circular object with a radius of $0.5\mathrm{m}$ and is assumed to be able to move in any direction on a planar surface, with the state at time step $t$, $\mathbf{s}_{t}\in\mathbb{R}^{2}$, consisting of the position of the pedestrian in two dimensional space. In these simulations, we consider a single pedestrian interacting with an autonomous vehicle. However, we note that our proposed algorithm applies to problems with multiple pedestrians, or problems with both pedestrians and other vehicles.
C.8 Synthetic Data
To train the neural network used in generating synthetic data, original pedestrian states were used as example outputs: $\mathbf{s}_{1:T}$ represents all of the states in one trajectory with $T$ time steps. Simultaneously, the corresponding inputs to each trajectory, $\mathbf{r}_{1:T}$, were generated by taking the starting and ending states of the trajectory so that $\mathbf{r}_{1}:=\mathbf{s}_{1}$ and $\mathbf{r}_{T}:=\mathbf{s}_{T}$ and then finding linear interpolations of these states. New, synthetic trajectories were generated after training by creating new rectilinear trajectories to serve as inputs to the network: as average pedestrian speeds and the angles between their starting and ending states each approximated a normal distribution, new rectilinear trajectories were created by sampling from normal distributions following the parameters matching those of the real data and passing the resulting trajectories through the network. The vertical direction then had a 50% chance of being inverted, and the horizontal starting position of the pedestrian relative to the starting position of the ego vehicle was normally distributed as the trajectories drawn from the VCI dataset were.
C.9 Pedestrian Behavior
For each simulation, the pedestrian adopts one of two modes of behavior for the whole trajectory. For each of the $10$ trajectories in the test set, the pedestrian either follows the selected trajectory throughout while crossing the crosswalk (representing in-distribution behavior for the predictor ensemble), referred to as nominal behavior, or the pedestrian can follow one of these trajectories for a set period of time (in these simulations, $1.3$ seconds) and then suddenly change their behavior by continuously moving towards the ego vehicle at maximum speed, referred to as insurance fraud behavior, representing a pedestrian attempting to force a collision with the ego vehicle. This presents both a worst-case scenario for the ego vehicle, which must prevent a collision despite the restrictions on its movements and a shift in distribution that the OOD detector must identify and respond to over the course of the simulation.
C.10 Statistical Analysis of Safety Assurances on Real and Synthetic Data
@@
Here, we analyze the statistical guarantees of our out-of-distribution detector empirically. We note that conformal prediction only provides probabilistic guarantees on the false-positive rate (i.e., on an incorrect identification of a data point in a nominal trajectory as an out-of-distribution sample). To assess the tightness of this guarantee, we apply the out-of-distribution detector at $27$ time steps in each of the $10$ nominal simulations, and at the $23$ corresponding time steps after the shift in pedestrian behavior from nominal to OOD behavior in each of the $10$ insurance fraud simulations. We summarize these results in Table 1, noting that the OOD detector has a false-positive rate of ${4.4\%}$, which is close to the probabilistic error rate of ${4\%}$ used as a calibration setpoint for conformal prediction. In Figure 7, we show a simulation where the detector incorrectly classifies a nominal behavior as being OOD. However, our detector is able to recover from incorrect classifications and subsequently classifies the behavior of the pedestrian as being in-distribution.
@@
Table 1: Confusion matrix showing the performance of the out-of-distribution (OOD) detector on the pedestrian trajectory data. The calibrated error rate $\delta\approx 4\%$ is close to the empirical false-positive rate (shown in bold).
@@
Figure 7: An instance of recovery from misclassification: in this simulation, the detector wrongly classifies nominal behavior as OOD, leading to a temporary switch from the nominal MPC controller (blue trajectory) to the conservative, provably-safe controller (red trajectory), before returning to the correct classification.
@@
Figure 8: Architecture of the LSTM model used to generate synthetic trajectory data for statistical analysis.
Figure 9: Beta distribution for calibration set of 100 samples and marginal probability of $96\%$, with an empirical coverage rate of $95.6\%$.
We note that the discrepancy between the observed error rate and the statistical guarantee provided by conformal prediction is expected: more specifically, since each of the trajectories are sampled i.i.d. from the probability distribution $\mathcal{D}$, the distribution of coverage for a threshold calibrated with $C=\rho^{(K)}$ follows a Beta distribution, with shape parameters $(a,b)=(N+1-K,K)$ (Angelopoulos and Bates, 2021). In Figure 9, we show the empirical distribution of coverage, in addition to the associated Beta distribution. We generate the empirical distribution using synthetic trajectories generated by an LSTM (described in Figure 9), trained on our training dataset, where the trajectories for the calibration and validation datasets are independently sampled without replacement. For each of $3000$ independent trials, $100$ independent trajectories are first generated and used to create a new calibration set of $100$ i.i.d. data points, with a new threshold set equal to the covariance measure associated with ${K=97}$ (the neural networks used to generate the covariance score are the same as during the simulations and were not retrained for this process). Subsequently, the covariance score for each of the data points in a further $150$ independently sampled trajectories is compared to the newly calibrated threshold, with the empirical coverage achieved during each trial being defined as the fraction of the data points in these $150$ trajectories with a covariance score less than or equal to the calibrated threshold. The number of trials that are associated with a given empirical coverage, out of the $3000$ in total, are proportional to the probability density described by the Beta distribution. Figure 9 shows that the empirical distribution closely matches the Beta distribution.
@@
C.11 Trajectron++, a Large-Scale Trajectory-Prediction Model
The Trajectron++ predictor is multi-modal, with the predicted positions of a given agent (which can be a pedestrian or a vehicle) represented as a Gaussian mixture model (GMM). At each time step, the system produces a set of 25 trajectories for each agent, with each of these trajectories representing a mode and having an assigned probability as part of a categorical distribution (there are several ways that these modes can be sampled from or combined to generate a single predicted trajectory, but this was not used over the course of our experiments). Each of the predicted trajectories corresponding to an individual mode consisted of 12 time steps, with each time step represented by the mean and covariance of a 2D Gaussian distribution.
The multi-modal nature of a model such as Trajectron++ can be leveraged to detect an increase in uncertainty, most notably through the entropy of the model as a whole or that of its constituent modes. In our case, uncertainty was estimated using the weighted sum of the determinants of the covariance matrices for the first predicted time step in each mode of the model (since the determinant of the covariance matrix has an affine relationship to the entropy of the individual Gaussian distribution):
$\rho_{t+1}=\sum_{z\in\mathcal{Z}}p_{\theta}(z|\mathbf{s})|\Sigma_{t+1}^{(z)}|,$ (9)
where $z$ is an integer representing a given mode within the model, $p(z|\mathbf{s})$ is the probability of this mode in the categorical distribution at the current time step, $\Sigma_{t+1}^{(z)}$ is the covariance matrix for the Gaussian distribution representing the first predicted time step for the mode and $\mathcal{Z}$ is the set of the integers representing all of the modes in the GMM. As these experiments are specifically intended to demonstrate outlier detection with this type of model and take place in cluttered environments with a large variety of interacting agents, the behavior of the ego vehicle follows that described in the original dataset, rather than being guided by MPC. The ego vehicle makes no attempt to dodge pedestrians displaying insurance fraud behavior.
Dataset. In order to demonstrate the use of conformal prediction to enable the calibration of out-of-distribution detection on a commercial-scale model in a realistic scenario, we evaluate the performance of the detection algorithm in scenarios included in the nuScenes (Caesar et al., 2020) dataset, which is taken from actual recorded vehicle and pedestrian movements in several drives on public roads, sampled at 2Hz. In the dataset, multiple pedestrians and other vehicles are visible simultaneously, and different agents appear in these scenes for different amounts of time. Each scene lasts 20 seconds, with an agent that is visible throughout the scene appearing for 41 time steps. Of the 850 annotated scenes available in the nuScenes dataset, which are assumed to be independent, 100 which contained adult pedestrians were removed and placed into a calibration set, with the trajectories used for training and testing drawn from the remaining 750 scenes. After training one instance of the Trajectron++ model, the trained model was used to perform calibration via conformal prediction, for the outlier detection: in each scene in the calibration set, one pedestrian is randomly selected, a time step in the corresponding trajectory is uniformly randomly selected and the uncertainty value is calculated for the prediction at this time step. This gives 100 independent data points used for the calibration procedure. Anomalous data was generated by modifying the behavior of adult pedestrians visible for 41 time steps, so that their positions would move, at each time step, towards the current position of the ego vehicle at the same speed as in the insurance fraud behavior in the previous experiments (all other data remained unmodified and no collision detection was considered). In total, 230 anomalous trajectories were analyzed in the test procedure, while there were a total of 6230 nominal trajectories of various lengths analyzed.
We repeated the statistical analysis on trajectories from the nuScenes dataset with Trajectron++, summarized in Table 2. In Table 2, the percentage values are computed relative to the ground-truth nominal and OOD behaviors: nominal cases cover 6230 trajectories of pedestrians in the dataset, while OOD cases use 230 trajectories with shifted pedestrian behavior. The calibrated error rate for individual data points is $\delta\approx 4\%$; a trajectory is classified as OOD when at least one data point in it has a nonconformity measure above the calibrated threshold and nominal otherwise.
Table 2: Confusion matrix showing the performance of the OOD detector through the various modes of the Trajectron++ predictor on trajectories from the nuScenes dataset.
True Behavior Identified as Nominal Identified as OOD Nominal $5361$ ($86.1\%$) $869$ ($13.9\%$) OOD $19$ ($8.3\%$) $211$ ($91.7\%$)
C.12 Statistical Analysis for Data Collection